# I Introduction

### v1

@@ II-A Cryo-CMOS Antenna Model
The dipole, illustrated in Fig. 2(b), comprises two co-linear copper arms deposited on a Silicon Dioxide (SiO2) layer, where the initial resonant length can be estimated as , where denotes the free-space wavelength and represents the effective permittivity of the antenna, with being the permittivity of the substrate. The final optimal dimensions were refined through iterative adjustments of the length of the dipole’s arms.
@@ III Performance Evaluation
III-A On-Chip Antenna Inside the Cryostat
The integrated on-chip antenna inside the cryostat is simulated using two discrete ports placed on GSSG pads, excited by impulse signals. The antenna design is fine-tuned to achieve efficient radiation and proper impedance matching to the cryostat environment. The optimized dipole length has been determined to be mm. As shown in Fig. 2(c), the reflection coefficient of the transmitter antenna exhibits a resonance close to 28 GHz with a minimum value of -21 dB, indicating a proper impedance matching and enhanced radiation performance within the cryostat.
The spatial field distribution of the antenna is observed in Fig. 3. The antennas are placed 6 cm above the second cooling plate to minimize the effects of the structure. The figure shows how the energy is distributed primarily through the PCB that holds the antennas, within its same temperature level. Due to reflections and scattering effects, the signal energy also travels to other temperature levels, with some radiation arriving to the cooling plate below. Significant amount of radiation arrives to the antennas that are co-lateral, either because the differential feed radiates or because of common substrate effects or reflections from the tubes or enclosure present in the environment.
@@ III-B Wireless Channel
In Fig. 4(c), the RMS delay spread of the CIR is obtained as detailed in prior sections. It is observed that the delay spread is relatively high, between 0.3 and 0.7 ns, leading to coherence bandwidths of a few GHz at most. The delay spread initially increases with distance, indicating that the main direct ray for communication weakens while distant reflections gain importance. At higher distances, the delay spread is slightly reduced as all reflections seem to be equally important. This is a usual behavior in enclosed environments [11].
The received power at each antenna position from to is shown in Fig. 4(d), illustrating how much energy is radiated towards each receiver antenna. Such localized power concentrations could potentially affect neighboring QPUs if the radiated power exceeds the allowable thermal dissipation threshold within the cryostat. Even though the separation between antennas ranges between and depending on the receiver antenna position, the received power remains largely independent of the distance, showing only minor variations. This is mainly due to energy reverberation within the cryostat. The tradeoff, then, lies in the energy-delay correlation since, owing to this reverberant behavior, both high energy and delay spread can be expected.
In Fig. 4(e), the SNR is measured as detailed in prior sections, and represented for varying bandwidths and different thermal noise power models at a transmit power of –30 dBm. The thermal noise added at the receiver is inherently low at high frequencies and cryogenic temperatures due to low carrier mobility at low temperatures. Yet, when the bandwidth increases, the noise is proportionally increasing, thereby reducing the SNR. Furthermore, it is observed that there is a significant increase of SNR in 4 K compared with room temperature due to low noise figures.

### v2

@@ II-A Cryo-CMOS Antenna Model
The dipole, illustrated in Fig. 2(b), comprises two co-linear copper arms deposited on a Silicon Dioxide (SiO2) layer. Initially, the resonant length was estimated as , where denotes the free-space wavelength and represents the effective permittivity of the antenna, with being the permittivity of the substrate [26], [27]. The final optimal dimensions were refined through iterative adjustments of the length of the dipole’s arms.
@@ III Performance Evaluation
III-A On-Chip Antennas Inside the Cryostat
The integrated on-chip antenna inside the cryostat is simulated using two discrete ports placed on GSSG pads, excited by impulse signals. The antenna design is fine-tuned to achieve efficient radiation and 50 reference impedance to the cryostat environment [25]. The optimized dipole length has been determined to be mm. As shown in Fig. 2(c), the reflection coefficient of the transmitter antenna exhibits a resonance close to 28 GHz with a minimum value of -22 dB, indicating a proper impedance matching and enhanced radiation performance within the cryostat.
The spatial field distribution of the antenna is observed in Fig. 3. The antennas are placed on top of the second cooling plate to minimize the RF effects to the QPUs. The figure shows how the energy is distributed primarily through the PCB that holds the antennas, within its same temperature level. Due to reflections and scattering effects, the signal energy also travels to other temperature levels, with some radiation arriving to the cooling plate below. Significant amount of radiation arrives to the antennas that are co-lateral, either because the differential feed radiates or because of common substrate effects or reflections from the tubes or enclosure present in the environment.
@@ III-B Wireless Channel
In Fig. 4(c), the RMS delay spread of the CIR is obtained as detailed in prior sections. It is observed that the delay spread is relatively high, between 1.7 ns and 2.2 ns, leading to coherence bandwidths in MHz. The delay spread initially increases with distance, indicating that the main direct ray for communication weakens while distant reflections gain importance. At higher distances, the delay spread is slightly reduced as all reflections seem to be equally important. This is a usual behavior in enclosed environments [11].
The received power at each antenna position from to is shown in Fig. 4(d), illustrating how much energy is radiated towards each receiver antenna. Such localized power concentrations could potentially affect neighboring QPUs if the radiated power exceeds the allowable thermal dissipation threshold within the cryostat. As the separation between antennas ranges between and depending on the receiver antenna position, the received power varies despite of the distance. This is mainly due to energy reverberation within the cryostat. The tradeoff, then, lies in the energy-delay correlation since, owing to this reverberant behavior, both energy variation and delay spread can be expected.
In Fig. 4(e), the SNR is measured as detailed in Section II-C1 and represented for different thermal noise power models in varied bandwidths, at a transmit power of –30 dBm. The thermal noise added at the receiver is inherently low at high frequencies and cryogenic temperatures due to low carrier mobility at low temperatures. Yet, when the bandwidth increases, the noise is proportionally increasing, thereby reducing the SNR. Furthermore, it is observed that there is a significant increase of SNR in 4 K compared with room temperature due to low noise figures.